OpenStudy (anonymous):

Solve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. 1 )Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) 2)Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

Directrix (directrix):

I just set up #2 a few minutes ago. Let A = the number of adults going on the trip Let S = the number of students going on the trip Then, A + S = 150 Each adult ticket cost $45 so the total amount of money the adults spend on tickets is $45 times A. Each student ticket cost $28.50 so the total amount of money the students spend on tickets is $28.50 times S. Solve these two equations simultaneously: A + S = 150 45*A + 28.5*S = 4770 -------------------- http://openstudy.com/users/directrix#/updates/51db52a5e4b076f7da4001ab

OpenStudy (anonymous):

can you help me solve them

Directrix (directrix):

I'll do one and you do the other. You get to choose first. So, which one is yours?

OpenStudy (anonymous):

the second one

Directrix (directrix):

I guess this is mine, then. 1 )Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

Directrix (directrix):

Let T equal Tony's age Let B equal Belinda's age. T + B = 56 B = 8 + 2T ======= T + B = 56 T + (8 + 2T) = 56 3T + 8 = 56 3T = 48 T = 16 B = 56 - 16 = 40 Toney is 16 and Belinda is 40. Do check my work.

Directrix (directrix):

Your turn for #2.

OpenStudy (anonymous):

A + S = 150 45*A + 28.5*S = 4770 A+(28.5*a)=4770,

Directrix (directrix):

A + S = 150 --> so A = 150 - S 45 A + 28.5 S = 4770 --------------------- 45(150 - S) + 28.5 S = 4770 What is the next step?

Directrix (directrix):

45*150 - 45*S + 28.5*S = 4770 45*150 - 16.5*S = 4770 * means times Okay, what is 45 times 150 ?

OpenStudy (anonymous):

6,750

OpenStudy (texaschic101):

are you stuck ?

Directrix (directrix):

45*150 - 16.5*S = 4770 6,750 - 16.5S = 4770 -16.5 S = 4770 - 6,750 @celeste8o What is 4770 - 6,750 = ?

OpenStudy (texaschic101):

she could participate....this is not hard

OpenStudy (texaschic101):

you are giving her everything she needs

Directrix (directrix):

And, she's offline at that.

OpenStudy (texaschic101):

well...you tried